# A Survey on the Application of Large Language Models in Scenario-Based Testing of Automated Driving Systems

### v1

@@ IV-B5 LLM as an Executable Scenario Generator
Template filling refers to the process in which LLMs populate predefined scenario template with specific parameters in the to generate executable files. Within this category, the study presented in [44] is representative. In this work, Güzay et al. propose a method for generating SUMO scenario based on predefined templates. Structured prompts are designed to guide GPT-4 in producing XML files according to user-defined variables, such as the number of intersections and the length of road segments.
@@
End-to-end generation refers to the direct mapping of natural language or multimodal inputs into executable simulation files, without relying on predefined templates. This paradigm is first explored by Wang et al. [53], who prompt general-purpose LLMs to generate XOSC files. To enhance controllability, Miceli-Barone et al. [43] propose a dialogue-based system in which SCENIC programs are incrementally constructed through iterative user interactions. Extending this approach to the visual domain, Miao et al. [46] introduce ScriptGPT, which translates dashcam videos into SCENIC scripts using prompt-engineered video-language models. To address the issue of structural correctness, Elmaaroufi et al. [54] combine compositional prompting with compiler feedback to generate probabilistic scenarios from crash reports.
@@
Hybrid generation approach combines template-filling and end-to-end generation, enabling greater flexibility while maintaining scenario standardization. In [55], Zorin et al. integrate real-world traffic incident data scraped from websites with LLM-based scenario completion methods to bridge open-source intelligence and executable simulation file generation. Yang et al. [49] introduce SimCopilot, which translates natural language descriptions of object interactions into simulation-ready code using a language-to-interaction dataset, thereby balancing high-level abstraction with detailed execution. Lu et al. propose OmniTester [42] and AutoScenario [48], both of which incorporate prompt engineering, Retrieval-Augmented Generation (RAG), and external simulation tools (e.g., SUMO, CARLA) to enhance the realism, diversity, and controllability of generated scenarios. Similarly, Zhang et al. [56] propose ChatScene, which decomposes textual inputs into subcomponents and retrieves relevant domain-specific code snippets to generate safety-critical scenarios. Aiersilan et al. [47] present AutoSceneGen that leverages in-context learning to transform high-level textual descriptions into simulator-compatible configuration files, enabling batch generation of safety-critical scenarios in CARLA.
IV-C Scenario Selection
@@
Scenario selection is performed to efficiently filter concrete scenarios from a database, with the aim of identifying cases that may reveal potential issues in ADSs [19]. As a recently explored approach, realism assessment emphasizes evaluating the degree to which generated scenarios align with real-world conditions, thereby allowing only realistic cases to be selected in order to improve test validity. LLMs have been applied to support this evaluation process.
@@
Two representative studies have investigated the use of realism assessment for scenario selection. In the study by Wu et al. [51], LLMs are employed to evaluate the consistency of driving trajectories with real-world conditions. This is accomplished by applying standardized prompts to assess perturbed variants of the DeepScenario dataset [121], resulting in a robustness score that quantifies scenario realism. In a related effort, Fu et al. [52] combine a diffusion model with a VLM to generate, narrate, and interpret realistic driving videos. The realism of the generated outputs is verified to enhance scene understanding. Experiments conducted on the Waymo Open Dataset [122] demonstrate the effectiveness of this approach in advancing VLM applications for AD.
IV-D Test Execution
@@
In scenario-based testing, concrete scenarios are executed within test environments such as real-world roads, close-track, or simulation platforms, each supporting varying levels of X-in-the-loop (XiL) integration [123, 19]. The studies reviewed in this work primarily focus on the use of LLMs in fully simulated environment. In these contexts, LLMs are employed to dynamically adjust testing parameters, thereby enabling self-adaptive testing procedures.
IV-D1 Anomaly Detection
@@
ADSs are susceptible to failures resulting from system-level deficiencies, which are commonly referred to as anomalies. Detecting such anomalies requires advanced reasoning capabilities. In this context, LLMs have been utilized to identify perception system anomalies in real time during test execution phase. Elhafsi et al. [50] apply OWL-ViT [124], a vision transformer model designed for object-level visual understanding, to extract visual feature, and GPT-3.5 for reasoning to detect anomalies in ADSs within the CARLA simulation environment. Their approach outperforms traditional Out-Of-Distribution (OOD) detection methods while offering interpretable analysis, although limitations remain due to the reliability of object detection and LLM inference.
IV-D2 Simulation Setup Automation
The configuration of simulation environments remains a complex and expertise-intensive task. Recent advancements in LLMs offer a promising solution for automating simulation setup by translating natural language inputs into executable configurations, a direction explored by several recent studies. In [39, 47] and [48], LLMs are employed to modify traffic environments, generate traffic scenarios, and configure simulation files for platforms SUMO [39] and CARLA [47, 48]. These approaches support real-time adjustments, enhance scenario diversity, and enable end-to-end automation without manual intervention. Building upon this research direction, Yang et al. [49] construct virtual traffic scenes in the LGSVL simulator based on natural language descriptions and introduce a language-to-interaction dataset introduced to support benchmarking and further development in simulation setup automation.
IV-D3 Scenario Optimization
@@
During the execution of concrete scenarios in physical simulators, error messages may occur due to syntax errors in scenario definitions. Traditionally, the correction of such errors has been performed manually and offline. To improve efficiency, Lu et al. [42] utilize a LLM integrated with a RAG module to iteratively validate and refine scenarios based on simulation feedback from CARLA and SUMO. Similarly, in [43], an LLM is used to automatically correct and update SCENIC code in response to both simulation results and user feedback, until successful execution is achieved in CARLA. In [44], GPT-4 is applied to parse error messages generated by SUMO and to modify the simulation files accordingly. Beyond syntax correction, LLMs have also been utilized for generating and refining closed-loop control code. In [45], an LLM is leveraged to generate and iteratively refine control code from natural language descriptions based on simulation feedback in Esmini [125], thereby forming a closed-loop workflow integrating code generation, simulation, and correction.
@@
Furthermore, if simulated scenarios are not consistent with real-world driving data, the resulting simulations may lack reliability. To refine scenario fidelity during execution, Miao et al. [46] utilize LLM to perform similarity analysis between simulated scenarios and dashcam crash video. In case of mismatch, the LLM iteratively refines the SCENIC script by incorporating feedback from a similarity metric. Scenario elements are updated until the simulated scenario closely approximates the original crash event.
IV-E ADS Assessment
@@
In this phase, the performance of ADSs is quantified using key metrics, with Time-To-Collision (TTC) [126] being one widely adopted example. These statistical indicators are typically included in simulation test report. To automate this process, LLMs have been employed to generate simulation reports. In [39], a LLM is used to interpret SUMO output files, extract relevant metrics such as traffic density, travel time, and emissions, and synthesizes analytical summaries to support efficient evaluation and comparison of simulation outcomes.
Beyond simulation, LLMs have also been applied to the analysis of real-world ADS accidents. Xu et al. [40] explore the potential of LLMs to generate legal explanations for ADS-related accident cases. The limitations of the models are evaluated, and improvements are proposed through domain-specific adaption and enhanced contextual reasoning.
@@
In addition to safety assessment, LLMs have been explored as tools to evaluate the intelligence level of ADSs. You et al. [41] employ LLMs to perform hierarchical assessments by using CoT prompting combined with RAG. The models simulate human-like reasoning across multiple decision-making levels, and their outputs are validated through CARLA simulations and human evaluators.
V Model selection and adaption
@@
LLMs have begun to play a crucial role in scenario-based testing of ADSs, with numerous LLMs proposed from different organizations. Table I and Table II provide a comprehensive summary of the LLMs used in the reviewed studies, categorized according to various attributes.
USA | OpenAI | GPT-4 | - | ① | ❺ | STPA | [93] | - |
❹ | HARA | [91] | - | |||||
❸❹ | Visual scene editing | [100] | - | |||||
❷❹ | HARA | [92] | - | |||||
❹ | HARA | [95] | - | |||||
②⑤ | ❷❸❹ | Functional scenario generation | [47] | CARLA | ||||
❷❸❼ | Scenario generation | [42] | SUMO | |||||
❷❸❹❼ | DSL scenario generation | [79] | - | |||||
❸ | Scenario generation | [64] | - | |||||
AVUnit generation | [57] | LGSVL | ||||||
❹ | Accident report interpretation | [70] | LGSVL | |||||
② | AVUnit generation | [59] | LGSVL | |||||
Driving behavior generation | [77] | Esmini | ||||||
❷❹❼ | Functional scenario interpretation SCENIC generation | [56] | CARLA | |||||
❶❹ | XOSC generation | [55] | - | |||||
Cost function generation | [76] | - | ||||||
❶ | Scenario encoder generation | [82] | - | |||||
❷❸ | Multimodal data interpretation | [48] | CARLA | |||||
Scenario generation | SUMO | |||||||
❶❹❼ | Driving behavior generation | [84] | - | |||||
❶❸ | Functional scenario interpretation | [78] | - | |||||
Scenario generation | [66] | - | ||||||
❷❹ | Loss function generation | [81] | - | |||||
DSL scenario generation | [69] | LGSVL | ||||||
SCENIC generation | [43] | CARLA | ||||||
Traffic rule interpretation JSON generation | [68] | CARLA | ||||||
SUMO configuration generation | [44] | SUMO | ||||||
❷❸❹ | User language interpretation | [62] | Isaac Gym Blender | |||||
Driving trajectory generation | ||||||||
⑤ | ❶❹ | Code generation | [45] | Esmini | ||||
GPT-4o | - | ②⑤ | ❷❸❹ | Functional scenario generation | [47] | CARLA | ||
② | ❷ | SCENIC generation | [46] | CARLA | ||||
❷❸❹ | JSON generation | [58] | CARLA | |||||
❶❹ | Accident video classification | [67] | - | |||||
GPT-4o-mini | - | ② | ❸❹ | Scenario generation | [65] | CARLA | ||
❶❹ | Accident video classification | [67] | - | |||||
GPT-4v | - | ① | ❹❻ | Visual grounding | [89] | - | ||
Text-davinci-003 | - | ⑤ | ❶ | Semantic anomaly detection | [50] | CARLA | ||
CLIP [127] | ① | ❹ | Image retrieval | [87] | - | |||
❶❹ | Scenario retrieval | [86] | - | |||||
GPT-3 | - | ① | ❹ | STPA | [94] | - | ||
④ | ❹ | Scenario realism assessment | [51] | |||||
GPT-3.5 | - | ② | ❶❷❸❹❼ | SCENIC generation | [54] | CARLA | ||
❶❷ | Domain knowledge distillation | [71] | - | |||||
❻ | XOSC generation | [53] | - | |||||
❺ | Trajectory generation | [83] | - | |||||
① | ❷❸ | Trajectory generation | [97] | - | ||||
Text-Embeeding -Ada-002 | - | ① | ❸❹ | Visual scene editing | [100] | - | ||
Anthropic | Claude-instant-1.2 | - | ② | ❶❷❸❹❼ | SCENIC generation | [54] | CARLA | |
LMSYS | Vicuna-7B-1.5 [128] | ① | ❶❹ | Driving video generation | [99] | - |
UW Madison | LLaVA-v1.5-7b [129] | ① | ❷ | Driving log retrieval | [85] | - | ||
LLaVA-v1.5-13b [129] | ② | ❹❻ | BEV interpretation | [73] | - | |||
BakLLaVA [130] | ① | ❶❹ | Scenario retrieval | [86] | - | |||
ByteDance | LLaVA-NeXT- Interleave [131] | ② | ❶❹ | Accident video classification Visual grounding | [67] | - | ||
LLaMA [132] | ②⑤ | ❻ | SCENIC generation | [49] | LGSVL | |||
LLaMA-2 [133] | ① | ❻❽ | Scenario retrieval | [88] | LGSVL | |||
④ | ❹ | Scenario realism assessment | [51] | - | ||||
LLaMA-2-7B [133] | ② | ❻ | XOSC generation | [53] | - | |||
USA | Meta | LLaMA 3.1 | ②⑤⑥ | ❹❼ | Road generation | [39] | SUMO | |
② | ❸❹ | Functional scenario generation | [72] | MetaDrive | ||||
LLaMA-3 [134] | ② | ❶❹ | XOSC generation | [55] | - | |||
LLaMA 3-8B | ② | ❹❻ | Driving behavior generation | [61] | - | |||
TinyLlama [135] | ① | ❻ | Trajectory interpretation | [96] | - | |||
CodeLlama [136] | ⑤ | ❶❹ | Code generation | [45] | Esmini | |||
Salesforce | BLIP-2 [137] | ① | ❶❹ | Scenario retrieval | [86] | - | ||
Sentence-T5 [138] | ② | ❷❹❼ | SCENIC generation | [56] | CARLA | |||
CodeGemma [139] | ② | ❷ | SCENIC generation | [45] | Esmini | |||
Gemma-7B [140] | ① | ❷ | Driving log retrieval | [85] | - | |||
Gemini 1.0 Ultra | - | ② | ❷❹❼ | DSL scenario generation | [80] | - | ||
China | Alibaba | Qwen-turbo | - | ② | ❶ | Driving behavior generation | [60] | CARLA SUMO |
Baidu | Yiyan | - | ② | ❻ | XOSC generation | [53] | - | |
Tsinghua Uni. | GLM-4-Air [141] | ② | ❸❹ | Scenario generation | [65] | CARLA | ||
DeepSeek | DeepSeek-Coder [142] | ⑤ | ❶❹ | Code generation | [45] | Esmini | ||
DeepSeek-R1 [143] | ⑤ | ❶❹ | Code generation | [45] | Esmini | |||
② | ❷❼ | Driving behavior analysis | [75] | - | ||||
DeepSeek-V3 | ② | ❷❼ | Driving behavior analysis | [75] | - | |||
France | Mistral AI | Mistral-7B-Instruct | ② | ❷❹ | Scenario generation | [74] | - | |
Mistral-7B | ④ | ❹ | Scenario realism assessment | [51] | - | |||
Mistral-7B | ④ | ❶❹ | Code generation | [45] | Esmini | |||
NA | NA | NA | NA | ① | ❺ | STPA | [90] | - |
NA | NA | ② | ❹ | XOSC generation | [119] | Esmini |
V-A LLM Origin
Figure 7 illustrates the distribution of LLMs applied in scenario-based testing of ADSs, categorized by organization and region. The majority of applied models originate from organizations based in the USA, with OpenAI leading by a substantial margin. In fact, the number of studies employing OpenAI models exceeds the combined total from all other organizations. China ranks second, with five models developed by three organizations, including recent contributions from DeepSeek. Within Europe, France is the only contributing country, represented by models primarily released by Mistral AI. These findings are further substantiated by the “Country” and “Organization” columns in Table I, and Table II.

### v2

@@ III-B5 LLM as an Executable Scenario Generator
Template filling refers to the process in which LLMs populate a predefined scenario template with specific parameters to generate executable files. Within this category, the study presented in [Guzay.2023] is representative. In this work, Güzay et al. propose a method for generating SUMO test scenario based on predefined templates. Structured prompts are designed to guide GPT-4 in producing XML files according to user-defined variables, such as the number of intersections and the length of road segments. Results show executable simulations, though limitations arise from outdated model knowledge, inconsistent file outputs for complex prompts, and occasional structural errors like dead-end roads or misplaced vehicles.
@@
End-to-end generation refers to the direct mapping of natural language or multimodal inputs into executable simulation files, without relying on predefined templates. This paradigm is first explored by Wang et al. [Wang.2023], who prompt and fine-tune general-purpose LLMs to generate XOSC; executable files can be produced, but performance is unstable To enhance controllability, Miceli-Barone et al. [MiceliBarone.2023] incrementally construct SCENIC programs via dialogue; tests on 16 scenarios reach a 50% success rate, with risks of evaluation bias and persistent semantic inaccuracies in scenario details. Extending to the visual domain, Miao et al. [Miao.2024] introduce ScriptGPT, which translates dashcam videos into SCENIC scripts using prompt-engineered video-language models; scenario creation is reduced from hours to minutes and 64% of cases are fully automated, but performance drops in low-visibility scenes and relies on manually selected exemplars in prompt design. To improve structural correctness, Elmaaroufi et al. [elmaaroufi2024scenicnl] combine compositional prompting with compiler feedback to generate probabilistic scenarios from crash reports. On California AV reports, they achieve high syntactic correctness with generation times reduced to two minutes, though limited by reasoning latency, and difficulty with complex maneuvers.
Nevertheless, end-to-end generation has inherent limitations. Its performance is highly dependent on the input modality, which often provides only a partial view of the real world. For instance, dashcam-based approaches [Miao.2024] are constrained by the single-camera field of view, omitting context outside the frame. Likewise, natural language and crash reports typically leave many assumptions implicit, necessitating additional user clarification [MiceliBarone.2023] or probabilistic modeling necessary to capture uncertainty [elmaaroufi2024scenicnl]. Moreover, directly producing executable simulation files may lead to syntactic or semantic errors, which require compiler feedback and refinement to ensure structural correctness [Wang.2023]. These limitations indicate that while end-to-end approaches offer efficiency and automation, their reliability for safety-critical ADS testing still demands cautious evaluation.
@@
The hybrid generation approach combines template-filling and end-to-end generation, enabling greater flexibility while maintaining scenario standardization. In [Zorin.2024], Zorin et al. integrate real-world traffic incident data scraped from websites with LLM-based scenario completion methods to bridge open-source intelligence and executable simulation file generation. Yang et al. [Yang.2023] introduce SimCopilot, which translates natural language descriptions of object interactions into simulation-ready code using a language-to-interaction dataset, thereby balancing high-level abstraction with detailed execution. Lu et al. propose OmniTester [Lu.2024d] and AutoScenario [Lu.2024], both of which incorporate prompt engineering, RAG, and external simulation tools (e.g., SUMO, CARLA) to enhance the realism, diversity, and controllability of generated scenarios. Similarly, Zhang et al. [Zhang.2024b] propose ChatScene, which decomposes textual inputs into subcomponents and retrieves relevant domain-specific code snippets to generate safety-critical scenarios. Aiersilan et al. [aasi2024generating] present AutoSceneGen that leverages in-context learning to transform high-level textual descriptions into simulator-compatible configuration files, enabling batch generation of safety-critical scenarios in CARLA.
In addition to text-based pipelines, hybrid approaches have also been extended to multimodal inputs. Ruan et al. [Ruan.2024] use LLMs to decompose NL descriptions into structured scenario elements, retrieving relevant road segments and planning multi-agent behaviors before rendering in CARLA. Meanwhile, Tian et al. [tian2025lmmenhancedsafetycriticalscenariogeneration] utilize LMMs to extract abstract scenarios from non-accident traffic videos; through multimodal few-shot CoT prompting, these are converted into concrete test scenarios, and a two-layer optimization quantifies the behavioral difference between the vehicle under test and human drivers, thereby exposing safety violation of ADSs.
III-B6 Comparative Analysis with Traditional Approaches
Traditional scenario generation follows three paradigms [ding2023survey]: data-driven methods learn distributions from collected datasets, knowledge-driven methods encode experts’ domain knowledge, and adversarial generation methods optimize parameters to maximize system failures. LLM-driven approaches introduce a fourth paradigm that converts NL descriptions into structured scenarios via reasoning. This shift from parameter optimization to semantic reasoning distinguishes LLMs from traditional methods but requires empirical validation.
Performance Comparison in Urban Scenarios
Table II presents quantitative comparison across four methodological paradigms, evaluated on SafeBench [xu2022safebench], a unified benchmarking platform for large-scale safety evaluation of AD algorithms. The metrics are derived from eight representative scenarios to ensure fair comparison. Three metrics are employed following definitions in [xu2022safebench]: CR for safety assessment, OS as a composite metric aggregating all SafeBench evaluation dimensions, and ADE measuring scenario diversity through trajectory variation. Compared methods include DiffScene [xu2025diffscene], CS [scenario_runner2025], AT [zhang2022adversarial], LC [ding2020learning], AS [wang2021advsim], SCENEGE [liu2025adversarial], ChatScene [Zhang.2024b], and T2T [sheng2025talk2traffic].
Across the eight urban scenarios, LLM‑driven approaches consistently lead on safety‑effectiveness (higher CR, lower OS) and deliver the greatest diversity (higher ADE), outperforming knowledge‑ and adversarial‑based baselines. This advantage likely stems from LLMs’ capability to reason about semantic scene composition and contextual relationships, rather than relying solely on parameter optimization or fixed templates. Data‑driven results are sparse on SafeBench, limiting comparability, but the overall trend supports the claim that LLMs strengthen both robustness and diversity in scenario‑based testing.
Comparing ChatScene [Zhang.2024b] and DiffScene [xu2025diffscene] illustrates this paradigm shift: while the former leverages LLM reasoning to generate context-aware and safety-critical scenarios, the latter relies on diffusion-based statistical learning without semantic interpretability. This highlights how LLM-driven approaches improve semantic diversity and controllability in scenario-based testing, yet their performance still depends on prompt quality and domain-specific fine-tuning, constraining reproducibility across datasets.
| Metric | Cat. | Method | Scenarios | Avg. | |||||||
| Straight Obstacle | Turning Obstacle | Lane Changing | Vehicle Passing | Red-light Running | Unprotected Left-turn | Right- turn | Crossing Negotiation | ||||
[…]
| ② | CS | 0.45 | 0.61 | 0.89 | 0.87 | 0.63 | 0.69 | 0.68 | 0.60 | 0.676 | |
| AT | 0.50 | 0.31 | 0.78 | 0.82 | 0.71 | 0.68 | 0.59 | 0.62 | 0.627 | ||
[…]
| ④ | ChatScene | 0.89 | 0.70 | 0.95 | 0.93 | 0.79 | 0.75 | 0.78 | 0.86 | 0.831 | |
| T2T | 0.913 | 0.780 | 0.893 | 0.947 | 0.900 | 0.833 | 0.860 | 0.893 | 0.877 | ||
[…]
| ② | CS | 0.698 | 0.567 | 0.489 | 0.490 | 0.641 | 0.613 | 0.505 | 0.579 | 0.573 | |
| AT | 0.668 | 0.714 | 0.538 | 0.505 | 0.607 | 0.620 | 0.545 | 0.569 | 0.596 | ||
[…]
| ④ | ChatScene | 0.470 | 0.522 | 0.434 | 0.440 | 0.537 | 0.560 | 0.474 | 0.421 | 0.482 | |
| T2T | 0.475 | 0.481 | 0.461 | 0.437 | 0.496 | 0.539 | 0.449 | 0.422 | 0.471 | ||
[…]
| ② | CS | 0.348 | 1.668 | 0.410 | 0.224 | 0.338 | 0.385 | 0.299 | 0.507 | 0.507 | |
| AT | 0.683 | 1.236 | 3.762 | 1.001 | 0.931 | 1.720 | 1.921 | 2.301 | 1.694 | ||
[…]
| ④ | ChatScene | 4.398 | 4.063 | 5.706 | 7.383 | 3.848 | 3.740 | 3.613 | 3.784 | 4.567 | |
| T2T | - | - | - | - | - | - | - | - | - | ||
Performance Comparison in Rural and Highway Scenarios
Few studies focus on rural scenario generation. Traditional approaches mainly generate scenario on two-lane, two-way rural roads [VIGNE2024104800, oh2024exploring, branzi2021combined]. To the best of the authors’ knowledge, no LLM-driven approaches are available; a comparison is therefore not applicable. In highway settings, traditional methods are predominantly data-driven, based on the authors’ survey [knies2020data, wei2025data, muslim2023cut, jing2024efficient, yang2023scegan, van2023deep]. LLM-driven approaches have also been explored [10588843, Chang.2024, tian2024enhancing]. However, diverse metrics across studies precluded fair cross-study comparison. This lack of methodological consistency across existing studies limits the interpretability and benchmarking value of current findings, highlighting the need for standardized evaluation protocols for LLM-based scenario generation.
III-C Scenario Selection
@@
Scenario selection is performed to efficiently filter concrete scenarios from a database, with the aim of identifying cases that may reveal potential issues in ADSs [Riedmaier.2020]. As a recently explored approach, realism assessment emphasizes evaluating the degree to which generated scenarios align with real-world conditions, thereby allowing only realistic cases to be selected in order to improve test validity. LLMs have been applied to support this evaluation process.
@@
Two representative studies have investigated the use of realism assessment for scenario selection. In the study by Wu et al. [Wu.2024], LLMs are employed to evaluate the consistency of driving trajectories with real-world conditions. This is accomplished by applying standardized prompts to assess perturbed variants of the DeepScenario dataset [lu2023deepscenario], resulting in a robustness score that quantifies scenario realism. In a related effort, Fu et al. [fu2024drivegenvlm] combine a diffusion model with a VLM to generate, narrate, and interpret realistic driving videos. The realism of the generated outputs is verified to enhance scene understanding. Experiments conducted on the WOD [sun2020scalability] demonstrate the effectiveness of this approach in advancing VLM applications for AD.
Beyond realism assessment, ensuring alignment between selected scenarios and the ADS’s ODD is equally crucial. However, ODD specifications are often expressed through semantic labels that pose challenges for automated processing. To the best of the authors’ knowledge, no studies have employed LLMs for ODD-scenario matching in the selection phase. A potential solution could adapt the approach proposed by Hildebrandt et al. [hildebrandt2024odd], which reformulates ODD semantic dimensions into yes-or-no queries to evaluate sensor data compliance, enabling similar adherence assessment for scenario selection.
III-D Test Execution
@@
In scenario-based testing, concrete test scenarios are executed within test environments such as real-world roads, close-track, or simulation platforms, each supporting varying levels of XiL integration [thorn2018framework, Riedmaier.2020]. The studies reviewed in this work primarily focus on the use of LLMs in fully simulated environments. In these contexts, LLMs are employed to dynamically adjust testing parameters, thereby enabling self-adaptive testing procedures.
III-D1 Anomaly Detection
@@
ADSs are susceptible to failures resulting from system-level deficiencies, which are commonly referred to as anomalies. Detecting such anomalies requires advanced reasoning capabilities. In this context, LLMs have been utilized to identify perception system anomalies in real time during the test execution phase. Elhafsi et al. [Elhafsi.2023] apply OWL-ViT [minderer2022simple], a vision transformer model, to parse camera images from CARLA into object-level scene descriptions, which are subsequently analyzed by GPT-3.5 to reason about potential anomalies in ADSs. Their approach reports improvements over OOD baselines while offering interpretable analysis, although performance is bounded by upstream detector errors and LLM inference variability.
III-D2 Simulation Setup Automation
The configuration of simulation environments remains a complex and expertise-intensive task. Recent studies leverage LLMs to streamline this process in two ways. One approach focuses on parameter specification, where LLMs generate structured parameters to guide simulator modules (e.g., ChatSUMO [Li.2024]). The other targets configuration generation, where LLMs directly produce executable scripts or configuration files (e.g., AutoSceneGen [aasi2024generating], AutoScenario [Lu.2024], and SimCopilot [Yang.2023]).
ChatSUMO [Li.2024] exemplifies the parameter-specification approach by enabling natural-language-based traffic flow generation and modification in SUMO. Validation shows high accuracy in road network generation, reduced setup time across both abstract and real-world networks, and user studies confirming improved usability. However, its reliance on potentially outdated OpenStreetMap666https://www.openstreetmap.org data and its inability to adapt to real-time events such as accidents or temporary road closures remain key limitations.
AutoSceneGen [aasi2024generating] converts user-defined text into CARLA-executable scripts via in-context learning. Experiments show that motion planners trained on its generated data outperform those trained solely on real data, though post-editing and limited explainability persist. AutoScenario [Lu.2024] leverages multimodal inputs (text, images, and videos) and LLM reasoning to generate scenarios in SUMO and CARLA, validated using conformity metrics and diversity metrics, but its lack of photorealism constrains fidelity. SimCopilot [Yang.2023] translates NL descriptions of object interactions into LGSVL-executable code, demonstrating strong performance across 120,000 interaction descriptions in six road topologies, though challenges remain in motion translation accuracy, handling of multi-object interactions, and generalization without off-road or traffic-violating behaviors.
III-D3 Scenario Optimization
@@
During the execution of concrete scenarios in physical simulators, error messages may occur due to syntax errors in scenario definitions. Traditionally, the correction of such errors has been performed manually and offline. To improve efficiency, Lu et al. [Lu.2024d] utilize a LLM integrated with a RAG module to iteratively validate and refine scenarios based on simulation feedback from CARLA and SUMO. Similarly, in [MiceliBarone.2023], an LLM is used to automatically correct and update SCENIC code in response to both simulation results and user feedback, until successful execution is achieved in CARLA. In [Guzay.2023], GPT-4 is applied to parse error messages generated by SUMO and to modify the simulation files accordingly. Beyond syntax correction, LLMs have also been utilized for generating and refining closed-loop control code. In [nouri2025simulation], an LLM is leveraged to generate and iteratively refine control code from natural language descriptions based on simulation feedback in Esmini [esmini], thereby forming a closed-loop workflow integrating code generation, simulation, and correction.
@@
Furthermore, if simulated scenarios are not consistent with real-world driving data, the resulting simulations may lack reliability. To refine scenario fidelity during execution, Miao et al. [Miao.2024] utilize LLMs to perform similarity analysis between simulated scenarios and dashcam crash video. Specifically, FeatureGPT, a GPT-4o-based video language model, estimates the probabilities of predefined features (e.g., weather, road type, vehicle maneuvers) in both sources and compares them against threshold values to detect inconsistencies. In case of mismatch, the LLM iteratively refines the SCENIC script by incorporating feedback from a similarity metric. Scenario elements are updated until the simulated scenario closely approximates the original crash event.
III-E ADS Assessment
@@
In this phase, the performance of ADSs is quantified using key metrics, with TTC [ozbay2008derivation] being one widely adopted surrogate safety metric [Westhofen2023]. This metric is often included in simulation test reports, though it represents only one approach to safety assessment. To automate this process, LLMs have been employed to generate simulation reports. In [Li.2024], a LLM is used to interpret SUMO output files, extract relevant metrics such as traffic density, travel time, and emissions, and synthesizes analytical summaries to support efficient evaluation and comparison of simulation outcomes.
Beyond simulation, LLMs have also been applied to the analysis of real-world ADS accidents. Xu et al. [Xu2024WhatTE] reveal the potential of LLMs to generate legal explanations for ADS-related accident cases through reviewing 31 papers. They find that while LLMs show promise in legal documents processing and consulting, they still falls short of meeting lawyer’s specific explanatory needs in ADS accident cases due to hallucination, limited domain-knowledge, and black-box reasoning mechanism.
@@
In addition to safety assessment, LLMs have been explored as tools to evaluate the intelligence level of ADSs. You et al. [you2025comprehensive] employ LLMs to perform hierarchical assessments by using CoT prompting combined with RAG. The models simulate human-like reasoning across multiple decision-making levels, and their outputs are validated through CARLA simulations and human evaluators.
IV Model selection and adaptation
@@
LLMs have begun to play a crucial role in scenario-based testing of ADSs, with numerous LLMs proposed from different organizations. Table IV and Table V provide a comprehensive summary of the LLMs used in the reviewed studies, categorized according to various attributes.
IV-A LLM Origin
Figure 9 illustrates the distribution of LLMs applied in scenario-based testing of ADSs, categorized by organization and region. The majority of applied models originate from organizations based in the USA, with OpenAI leading by a substantial margin. In fact, the number of studies employing OpenAI models exceeds the combined total from all other organizations. China ranks second, with five models developed by three organizations, including recent contributions from DeepSeek. Within Europe, France is the only contributing country, represented by models primarily released by Mistral AI. These findings are further substantiated by the “Country” and “Organization” columns in Table IV, and Table V.